OpenStudy (anonymous):

Find the slope of line passing through the points (5, 9) and (5, -2) (A)-11 (B)0 (C)11 (D)undefined

jigglypuff314 (jigglypuff314):

\[slope = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]but when you try plug in and solving it... numerator: -2 - 9 = -11 denominator: 5 - 5 = 0 you cannot have a 0 in the denominator therefore the slope is undefined

OpenStudy (anonymous):

thanks!! can u plz help me with one more???

jigglypuff314 (jigglypuff314):

sure :)

OpenStudy (anonymous):

thanks :) Find the slope of line passing through the points (4,8) and (6, -2).

OpenStudy (anonymous):

1/5 5 -1/5 -5

jigglypuff314 (jigglypuff314):

you can use the same slope form as before :) \[slope = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]so in numerator: -2 - 8 = -10 in denominator: 6 - 4 = 2 so you get -10/2 =

OpenStudy (anonymous):

so its -1/5 right? sry my brain is fried lolz

jigglypuff314 (jigglypuff314):

negative 10 divided by 2 equals negative 5

OpenStudy (anonymous):

O!!! ok i get it..oops lolz XD thank you so much!!! u r AWESOME!!!

jigglypuff314 (jigglypuff314):

you're welcome :)